OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). (explain help please)

OpenStudy (tamtoan):

what is the slope of the given line? can you tell?

OpenStudy (anonymous):

Parallel lines have the same slope. Find the slope of the given line and use the point given to create an equation

OpenStudy (anonymous):

your line is 3x -y =-3 => y = 3x +3 => y = 3 ( x + 1 ) this is your line that looks like this y = m ( x - x0 ) so { m = 3 } that is the SLOPE of the line to get the PARALLEL line on that Line it would be { mP = -1/3 } because ( m * mP ) sould be ( -1 ) the SLOPE of a Line * SLOPE of PARALLEL Line on it = -1

OpenStudy (anonymous):

( y - y0 ) = m ( x - x0 ) mP= - 1/3 A( 4 , -2 ) x0 y0 So the equation Would Be ( y + 2 ) = -1/3 ( x - 4 )

OpenStudy (tamtoan):

sofwon, isn't mP is slope of perpendicular line? :) slope of parallel lines should be the same

OpenStudy (anonymous):

oh soooooooooo soooooooooory got mixed up so it would be the same slope mP=3 y - y0 = m ( x - x0 ) y + 2 = 3 (x - 4)